OpenStudy (anonymous):

Solve the following system of equations using the substitution method. F = C F = 56 + (9/5)C

OpenStudy (anonymous):

F=56+(9/5)F F-9F/5=56 (5F-9F)/5=56 -4F=56*5 F=280/-4 F=-70 F=-70,C=-70